# Policymaker meetings as heteroscedasticity shifters:

### v1

@@ Abstract
We propose a novel approach to identification in structural vector autoregressions (SVARs) that uses external instruments for heteroscedasticiy of a structural shock of interest. This approach does not require lead/lag exogeneity for identification, does not require heteroskedasticity to be persistent, and facilitates interpretation of the structural shocks. To implement this identification approach in applications, we develop a new method for simultaneous inference of structural impulse responses and other parameters, employing a dependent wild-bootstrap of local projection estimators. This method is robust to an arbitrary number of unit roots and cointegration relationships (including seasonal unit roots), time-varying local means and drifts, and conditional heteroskedasticity of unknown form and can be used with other identification schemes, including Cholesky and the conventional external IV.
We show how to construct pointwise and simultaneous confidence bounds for structural impulse responses and how to compute smoothed local projections with the corresponding confidence bounds. Using simulated data from a standard log-linearized DSGE model, we show that the method can reliably recover the true impulse responses in realistic datasets. As an empirical application, we adopt the proposed method in order to identify monetary policy shock using the dates of Federal Open Market Committee (FOMC) meetings in a standard six-variable VAR. The robustness of our identification and inference methods allows us to construct an instrumental variable for monetary policy shock that dates back to 1965. The resulting impulse response functions (IRFs) for all variables align with the classical Cholesky identification scheme and are different from the narative sign restricted Bayesian VAR estimates. In particular, the response to inflation manifests a price puzzle that is indicative of the cost channel of the interest rates.
Keywords: SVAR, conditional heteroscedasticity, unit roots, cointegration, identification, simultaneous inference, smoothed local projections, narative sign restrictions, minimum distance estimator, monetary policy shocks.
@@ 1 Introduction
Unexpected actions by policymakers provide a valuable source of variation that helps to identify how the economy would react to future policy actions (Sims, 1980). Unfortunately, it is often difficult to detect the unexpected component of the actions using the publicly available data. The true plan of future actions and commitments of the policy makers can be kept secret from the public, and the pure unexpected part of the available policy announcements is always a matter of modeler’s assumptions. We propose an identification approach that remains agnostic about the actual size and directions of change in the policy variable (for example, interest rate) due to policymakers’ discretion. Instead, we only assume that the conditional variance of the unexpected change in the policy variable is higher during policymakers’ meetings than at other times. We show that the information contained in the meetings dates is sufficient to uniquely identify both the size and direction of impulse responses to the structural policy shock in a vector autoregression model (VAR).
@@
Our paper applies this novel approach to identification of the monetary policy shocks using the number of the US Federal Reserve Open Market Committee (FOMC) meetings per month as an instrument for heteroskedasticity of a monetary policy shock in a standard six-variable VAR model. Information available on the meetings dates goes back to at least 1965, which gives us instruments with sample sizes that were unprecedented in the monetary SVAR literature. Such long samples come with an additional set of complications. In particular, the inference procedure needs to be robust to the potential presence of time trends and unit roots, which makes the existing popular methods like residual-based bootstrap (Gonçalves and Kilian, 2004) inapplicable. Moreover, the use of the external instruments for the conditional heteroscedasticity requires a novel procedure that would allow joint inference on all the relevant parameters in the model.
@@
We address this challenge by proposing a novel inference procedure for the structural impulse responses (IRF) based on local projection (LP) estimators of Jordà (2005). Specifically, we propose both a formula and a proof of joint asymptotic normality of the local projection estimator in the potential presence of trends, cointegration and any number of unit roots. Next, we extend the proof of consistency of the joint HAC covariance estimator of Andrews (1991) to accomodate nonstationary regressors. Finally, we propose a version of the dependent wild bootstrap procedure of Shao (2010) that can mimic the joint distribution of the local projection estimators across multiple horizons and the estimates of structural parameters. The only regularity conditions that we impose on the non-explosive VAR model is that the one-step ahead forecast errors, jointly with the instrumental variable, follow a stationary vector process with a mixing property and certain higher-order moments necessary for consistency of the HAC estimator. To our knowledge, this provides the first proof of validity of inference on the IRF estimators that allows the general case of unstable but non-explosive VAR models with a general form of stationary conditional heteroscedasticity without pretests. [^2] Our results build upon the long literature on the non-stationary time series including Sims et al. (1990); Tsay and Tiao (1990); De Jong and Davidson (2000); Inoue and Kilian (2002); Montiel Olea and Plagborg-Møller (2021).
@@
The proposed procedure consists of three main steps. First, we estimate reduced-form impulse response functions (i.e. responses to one-step ahead linear forecast errors $\hat{\eta}_{it}$ for each time series $i$). This step is computed using local projection regressions. Second, we estimate contemporaneous impact of the structural shock of interest on a time series $i$ as the sample covariance between pairwise products of one-step ahead forecast errors $\hat{\eta}_{1t}\hat{\eta}_{it}$ and an instrumental variable correlated with the heteroscedasticity of the structural shock of interest. (The time series with index $1$ plays a special role since it is assumed to have non-zero correlation of squared forecast errors with the instrumental variable.) Third, we compute structural IRF for other horizons as products of the reduced form IRF and vector the contemporaneous structural responses after appropriate normalization. The dependent wild-bootstrap is then applied to the relevant score functions.
@@
The novel simultaneous inference on local projections has multiple uses. First, as emphasized in Brüggemann et al. (2016), it is important to make joint draws of the vector of the IRF coefficients and the structural impact vector to obtain valid confidence limits. Second, the procedure allows for simultaneous confidence bounds that cover the entire IRF across all horizons and, for example, allows for inference on the largest and smallest values of IRF. Third, one can compute smoothed local projection estimators and the corresponding confidence bounds using the minimum distance approach to improve efficiency of the estimates.
@@
We illustrate the proposed method in two applications. First, we show that the identification strategy can recover the true IRF of a monetary policy shock in simulated data based on a medium-sized DSGE model of US economy of Smets and Wouters (2007). Then we apply the method to study the impact of monetary policy shocks in six-variable VAR model using the data set of Antolín-Díaz and Rubio-Ramírez (2018). The variables include Real GDP, GDP deflator, commodity price index, total reserves, non-borrowed reserves, and Federal funds rate. The sample period is from January 1965 to November 2007. Our primary instrumental variable for conditional heteroscedasticity of the monetary policy shocks is the number of FOMC meetings per month, including telephone conferences and unscheduled meetings. For reduced form estimates, we use local projection with polynomial time trend with power up to $k\in\{0,4\}$ and the number of lags $p\in\{12,24\}$. We report various confidence bands: point-wise, simultaneous sup-t, smoothed point-wise and smoothed simultaneous bands.
@@
Our main empirical findings can be summarized as follows: (i) the identification through the heteroscedasticity IV results in IRF that are identical to the classical Cholesky identification a la Christiano et al. (1999);(ii) smoothed LP-IRF are noticeably different from the conventional recursive VAR estimates at the long horizons; (iii) effects on real GDP are comparable with BVAR identified using the narrataive sign restriction effects (Antolín-Díaz and Rubio-Ramírez, 2018), while the effects on the monetary shocks on inflation are much weaker in the novel scheme.
@@
The literature on identification through hetersocedasticity has a long history (for example, Feenstra (1994); Rigobon (2003); Lewis (2021)). To our knowledge, we are the first to propose an IV-type procedure that assumes that the conditional heteroscedasticity is not-persistent (c.f. Rigobon, 2003; Lewis, 2021), or structural shocks are fully independent and non-Gaussian (c.f. Lütkepohl and Netšunajev, 2017). Moreover, we do not assume any particular parametric model of the innovations and the instruments. In contrast to statistical identification procedures, our approach does not require full identification of all shocks and gives an intuitive interpretation of the identified shock of interest.
@@
In this paper, we apply our identification method to monetary policy shocks, but the external heterscedasticity IV can be proposed in other applications. For example, Känzig (2021) shows that OPEC meetings have high-frequency impacts on oil prices that can be used to identify oil supply shocks; Janzen et al. (2018) show that cotton storage levels change the heteroskedasticity of precautionary demand shocks. So we believe that our approach can be used more broadly.
@@
Our inference procedure is based on the insight that both the reduced-form impulse response and covariance estimators involved in the structural identification can be represented as a mean of vector-valued stationary time series of cross-products of the $1$-step and $h$-step ahead forecast errors. This representation allows us to employ the dependent wild bootstrap of Shao (2010) to make draws from the joint asymptotic distribution of the estimators without explicitly computing the long-run variance-covariance matrix estimator. This novel inference procedure can be used for many other identification approaches that include Cholesky (Christiano et al., 1999), sign restrictions (Gafarov et al., 2016), and external IV (Stock, 2008; Mertens and Ravn, 2013).
@@
The paper is structured as follows. Section 2 discusses the setup and the novel identification idea. Section 3 provides the estimation and inference algorithms with the corresponding theoretical proofs. Section 4 discusses the DSGE experiment. Section 5 has the empirical application to US monetary policy. Section 6 concludes. All the proofs and additional figures are provided in Appendix.
2 Model and the identification
@@ 2.1 Motivating example
where observations $\eta_{t}\in\mathbb{R}^{n}$ and $\varepsilon_{t}$ is an i.i.d. zero-mean process with $\mathbb{E}\varepsilon_{t}\varepsilon_{t}^{\prime}=I_{n}$ that is refereed to as structural shocks. [^3] The simultaneous equations model with lags were first proposed and studied in Haavelmo (1943); Mann and Wald (1943b). Suppose that we are interested in measuring the impact of the structural shock $\varepsilon_{1t}$ on the vector of observables $\eta_{t}$, which is given by the first column of $B=(b_{\cdot 1},\dots,b_{\cdot n})$. We can estimate $\Sigma_{\eta}\triangleq\mathbb{E}\eta_{t}\eta_{t}^{\prime}$ using the sample analog and the solve the equation $\Sigma_{\eta}=BB^{\prime}$ for the first column of $B$. However, there is a continuum of solutions to this equation. Therefore, additional assumptions or data are required to uniquely select $B$. The purpose of the paper is to propose a novel way of identifying $b_{\cdot 1}$ using external instruments for the conditional heteroscedasticity process of $\varepsilon_{t}$ and develop the inference procedure. This identification problem occurs in the more general SVAR models, which we will study later in the paper.
Before we state the novel identification proposal in the general form, let us consider a simple motivating example.
@@
such that $\xi_{it}$ is i.i.d. sequence with $\mathbb{E}\xi_{t}=0$ and $\mathbb{E}\xi_{t}\xi^{\prime}_{t}=I_{2}$; the process $\sigma_{it}$ is a stationary ergodic process that is independent of $\xi_{it}$. The volatility process $\sigma^{2}_{1t}$ for the shock of interest may be correlated with some observable time series of $Z_{t}$. Suppose that this time series $Z_{t}$ only increases volatility of the first structural shock, but does not change volatility for other shocks. We can then use it as an instrument for identification of $b_{\cdot 1}$. In contrast to the conventional proxy-VAR identification, the instrument $Z_{t}$ may be independent of $\xi_{1t}$. Indeed, suppose that
$\mathbb{E}(Z_{t}-\mathbb{E}Z_{t})(\sigma^{2}_{it}-\mathbb{E}\sigma^{2}_{it})=\rho_{i}$ (3)
@@
$\eta^{2}_{1t}$ $=(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t})(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t}).$ (4)
@@
$\mathbb{E}\eta^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})$ $=b_{11}^{2}\mathbb{E}\varepsilon^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})=b_{11}^{2}\rho_{1}$ (6)
@@
The same logic applies if $Z_{t}$ is correlated with $\xi_{t}$ as long as $\rho_{1}\neq 0$, $\rho_{2}=0$ and $\mathbb{E}\xi_{1t}\xi_{2t}(Z_{t}-\mathbb{E}Z_{t})=0$. ∎
@@
This assumption implies that the heteroscedasticity of the innovations for this time series must be correlated with the instrument $Z_{t}$. In the monetary shock example, one can by definition assume that the impact $b_{11}$ is non-zero for federal funds rate ($i=1$). The instrument can only change the conditional variance of the shock of interest and cannot result in conditional correlation of the other structural shocks.
@@
A recent work Lewis (2021) showed that in the case when heterscedasticity of a structural shock is persistent, one can use internal instruments $Z_{t}=\eta^{2}_{t-1}$. However, the approach can be applied more generally if appropriate external instruments are available, as in our empirical application.
@@
where $y_{t}$ is $n$-dimensional time series, $V\mu_{t}$ is a deterministic polynomial time trend (or time-varying drift for integrated processes) with $\mu_{t}=(1,t,t^{2},\dots,t^{k})$ for some $k$; $A_{1},\dots,A_{p},V$ are $(n\times n)$ parameter matrices, $\eta_{t}$ are serially uncorrelated forecast errors (innovations). We will assume that $p$ is finite and known.
@@
    $\eta_{t}$ is a martingale difference sequence (MDS), i.e. $\forall$ $t$ $\mathbb{E}[\eta_{t}|\mathcal{F}_{t-1}]=0$
@@
    $\mathbb{E}||\eta_{t}||^{8+2\epsilon}<\infty$ and $\mathbb{E}|Z|^{4+\epsilon}<\infty$ for some $\epsilon>0$.
@@
Assumption 2.1 implies that innovations are white noise, which is required for the consistency of LS estimators in stationary VAR models. The martingale difference requirement with respect to the sigma algebra $\mathcal{F}_{t-1}=\sigma(Y_{t-1},Y_{t-2},\dots)$ is only necessary for interpretation of the impulse response coefficients that will be introduced later in this section. Assumption 2.2 allows for many stationary models of conditional heteroscedasticity. [^4] See, for example, Assumption 2.1.(i-iii) in Brüggemann et al. (2016). This assumption, for example, allows for asymmetric conditional heteroscedasticity models like the ones considered in Rabemananjara and Zakoian (1993). Such models are ruled out by popular residual wild bootsrap methods (Gonçalves and Kilian, 2004). We do not consider unconditional heteroscedasticity in innovations, since it would eliminate the need for structural identification. The 8-th moment condition imposed in our paper is sufficient for consistency of the HAC estimators, but a weaker 4-th moment condition would be sufficient for asymptotic normality of the LP estimators. Assumption 2.4 is standard in papers on structural identification of the shocks. Assumption 2.5 is required for asymptotic normality of $\hat{\gamma}$ and would be satisfied, for example, if $\eta_{t}$ is MDS with respect to a sigma field $\mathcal{F}_{t-1}$ that also includes $Z_{t}$ and past values of $\eta_{t}$ (for example, it is possible that the number of meetings in month $t$ is fully predictable based on the past economic conditions and the past number of meetings). This assumption is not required for consistency of $\hat{\gamma}$ and the structural IRF.
@@
$C_{h,i,j}=\mathbb{E}(y_{t+h}|\eta_{t}=e_{j},\mathcal{F}_{t-1})-\mathbb{E}(y_{t+h}|\eta_{t}=0,\mathcal{F}_{t-1})$ (13)
where $e_{j}\in R^{n}$ is vector of zeros with $1$ in $j-$th position. These coefficients appear as regression coefficients in $H$-step ahead autoregression representation of (12),
@@
Assumption 3 allows for arbitrary number of unit roots, including complex unit roots, and cointegrating relations between $y_{t}$. It only rules out explosive roots. In practical terms, it means that one can apply our inference procedure to local projections estimated with or without trend in levels without seasonal adjustment. It can be done without any pretests for trend stationarity, abnsency of cointegration, or unit roots. We allow both a polynomial trend as in Sims et al. (1990) and complex unit roots as in Tsay and Tiao (1990). Under this assumption, the least squares estimator for the coefficients in equations (12) and (14) is consistent and has well-defined distribution limits as shown in Sims et al. (1990); Tsay and Tiao (1990) under conditionally homosecdastic MDS innovations $\eta_{t}$. In contrast, we allow for more general white noise innovations with mixing by using strong laws of large numbers and functional central limit theorems for strong mixing sequences (De Jong and Davidson, 2000; Rio et al., 2017).
@@
The coefficients $C_{1},\dots,C_{H}$ can be estimated using local projection estimators of Jordà (2005). However, our empirical strategy uses conditional heteroskedasticity in the structural shocks in the dataset that has unit root time series with potential cointegration and time trends. Furthermore, structural impulse responses depend on both estimators of $\Sigma_{\eta}$ and $C_{1},\dots,C_{H}$ and on covariances with external instruments $Z_{t}$. Existing inference methods cannot take into account all of these features together. Specifically, there are no proposals for joint inference of $\Sigma_{\eta}$ and $C_{0},\dots,C_{H}$ without pretests for unit roots.  [^5] Brüggemann et al. (2016) provide an example where commonly used residual wild-bootstrap is invalid for joint inference on $C_{i}$ and $\Sigma_{\eta}$ even in the stationary case. Their proposed procedure, residual block bootstrap, solves the issue for stationary VARs. Jentsch and Lunsford (2022) then extends this residual block bootstrap approach to allow for external instrumental variables, also in the stationary case. Inoue and Kilian (2002); Mikusheva (2012) allow for joint inference on individual $C_{h}$ and $\Sigma_{\eta}$, but impose an i.i.d. assumption on the innovations.
Consider the following example that illustrates the distinction between the single coefficient and joint inference on impulse responses in non-stationary case.

### v2

@@ Abstract
Structural vector autoregressions are used to compute impulse response functions (IRF) for persistent data. Existing multiple-parameter inference requires cumbersome pretesting for unit roots, cointegration, and trends with subsequent stationarization. To avoid pretesting, we propose a novel dependent wild bootstrap procedure for simultaneous inference on IRF using local projections (LP) estimated in levels in possibly nonstationary and heteroscedastic SVARs. The bootstrap also allows efficient smoothing of LP estimates.
We study IRF to US monetary policy identified using FOMC meetings count as an instrument for heteroscedasticity of monetary shocks. We validate our method using DSGE model simulations and alternative SVAR methods.
Keywords: SVAR, external instruments, conditional heteroscedasticity, unit roots, cointegration, identification, simultaneous inference, smoothed local projections, narrative sign restrictions, minimum distance estimator, monetary policy shocks.
@@ 1 Introduction
Unexpected actions by policymakers provide a valuable source of variation that helps to identify how the economy would react to future policy actions (Sims, 1980). Unfortunately, it is often difficult to detect the unexpected component of the actions using publicly available data. The true plan of future actions and commitments of the policymakers can be kept secret from the public, and the pure unexpected part of the available policy announcements is always a matter of modelers’ assumptions. We propose an identification approach that remains agnostic about the actual size and direction of change in the policy variable (for example, interest rate) due to policymakers’ discretion. Instead, we only assume that the conditional variance of the unexpected change in the policy variable is higher during policymakers’ meetings than at other times. We show that the information contained in the meeting dates is sufficient to uniquely identify both the size and direction of impulse responses to the structural policy shock in a vector autoregression model (VAR).
@@
Our paper applies this novel approach to the identification of monetary policy shocks using the number of US Federal Reserve Open Market Committee (FOMC) meetings per month as an instrument for heteroscedasticity of a monetary policy shock in a standard six-variable VAR model. The information available on the meeting dates goes back to at least 1965, which gives us instruments with sample sizes that were unprecedented in the monetary SVAR literature. Such long samples come with an additional set of complications. [^2] According to (Ramey, 2016), the longer sample period, on the one hand, includes more variation in interest rate and thus advantageous for monetary policy shock identification; on the other hand, it is accompanied by extra challenges such as structural breaks, regime shifts, state-dependence, and sign-dependence. In particular, the inference procedure needs to be robust to the potential presence of time trends and unit roots, which makes the existing popular methods like the residual-based bootstrap (Gonçalves and Kilian, 2004) inapplicable. Moreover, the use of external instruments for conditional heteroscedasticity requires a novel procedure that allows joint inference on all the relevant parameters in the model.
@@
We address this challenge by proposing a novel inference procedure for structural impulse responses (IRF) based on local projection (LP) estimators of Jordà (2005). Specifically, we propose both a formula and a proof of joint asymptotic normality of the local projection estimator in the potential presence of trends, cointegration, any number of unit roots, and any weakly-dependent stationary conditionally heteroscedastic innovations. Next, we extend the proof of the consistency of the joint HAC covariance estimator of Andrews (1991) to accommodate non-stationary regressors. Finally, we propose a version of the dependent wild bootstrap procedure of Shao (2010) that can mimic the joint distribution of the local projection estimators across multiple horizons and the estimates of structural parameters. The only regularity conditions that we impose on the non-explosive VAR model is that the one-step-ahead forecast errors, jointly with the instrumental variable, follow a stationary vector process with a mixing property and existence of certain higher-order moments, which are necessary for consistency of the HAC estimator. To our knowledge, this provides the first proof of the validity of inference on the IRF estimators that allows for the general case of unstable but nonexplosive VAR models with a general form of stationary conditional heteroscedasticity without pretests. [^3] Our results are built upon the long literature on non-stationary time series including Sims et al. (1990); Tsay and Tiao (1990); De Jong and Davidson (2000); Inoue and Kilian (2002); Montiel Olea and Plagborg-Møller (2021).
@@
The proposed procedure consists of three main steps. First, we estimate reduced-form impulse response functions (i.e. responses to one-step-ahead linear forecast errors $\hat{\eta}_{it}$ for each time series $i$). This step is computed using local projection regressions. Second, we estimate the contemporaneous impact of the structural shock of interest on time series $i$ as the sample covariance between pairwise products of one-step-ahead forecast errors $\hat{\eta}_{1t}\hat{\eta}_{it}$ and an instrumental variable that is correlated with the heteroscedasticity of the structural shock of interest. (The time series with index $1$ plays a special role since it is assumed to have a non-zero correlation of squared forecast errors with the instrumental variable.) Third, we compute structural IRF for other horizons as products of the reduced-form IRF and vector of contemporaneous structural responses after appropriate normalization. The dependent wild bootstrap is then applied to the relevant score functions.
@@
The novel simultaneous inference on local projections has multiple uses. First, as emphasized in Brüggemann et al. (2016), it is important to make joint draws of the vector of the IRF coefficients and the structural impact vector to obtain valid confidence limits. Second, the procedure allows simultaneous confidence bounds that cover the entire IRF across all horizons and, for example, allows for inference on the largest and smallest values of IRF. Third, one can compute smoothed local projection estimators and the corresponding confidence bounds using the minimum distance approach to improve the efficiency of the estimates.
@@
We illustrate the proposed method in two applications. First, we show that the identification strategy can recover the true IRF of a monetary policy shock in simulated data based on a medium-sized DSGE model of the US economy of Smets and Wouters (2007). Then we apply the method to study the impact of monetary policy shocks in a six-variable VAR model using the data set of Antolín-Díaz and Rubio-Ramirez (2018). The variables include Real GDP, GDP deflator, commodity price index, total reserves, non-borrowed reserves, and the Federal funds rate. The sample period is January 1965 to November 2007. Our primary instrumental variable for conditional heteroscedasticity of the monetary policy shocks is the number of FOMC meetings per month, including telephone conferences and unscheduled meetings. For reduced-form estimates, we use local projection with polynomial time trends with power up to $k\in\{0,4\}$ and the number of lags $p\in\{12,24\}$. We report various confidence bands: point-wise, simultaneous sup-t, smoothed point-wise, and smoothed simultaneous bands.
@@
Our main empirical findings can be summarized as follows: (i) the identification through the heteroscedasticity IV results in IRFs that are nearly identical to the classical Cholesky identification a la Christiano et al. (1999); (ii) smoothed LP-IRFs are noticeably different from the conventional recursive VAR estimates at the long horizons; (iii) effects on real GDP are comparable with BVAR identified using the narrative sign restriction effects (Antolín-Díaz and Rubio-Ramirez, 2018), while the effects of the monetary shocks on inflation are much weaker in the novel scheme.
@@
The literature on identification through heteroscedasticity has a long history (for example, Feenstra (1994); Rigobon (2003); Lewis (2021); Brunnermeier et al. (2021)). To our knowledge, we are the first to propose an IV-type procedure that assumes that the conditional heteroscedasticity is not persistent, or structural shocks are fully independent and non-Gaussian (c.f. Lütkepohl and Netšunajev, 2017). Moreover, we do not assume any particular parametric model of the innovations and the instruments. In contrast to statistical identification procedures, our approach does not require full identification of all shocks and gives an intuitive interpretation of the identified shock of interest.
@@
In this paper, we apply our identification method to monetary policy shocks, but the heteroscedasticity-based external IV can be proposed in other applications. [^4] There is a long-standing literature devoted to the identification of monetary policy shocks, including but not limited to Christiano et al. (1999), Romer and Romer (2004), Gürkaynak et al. (2005), Gertler and Karadi (2015), Stock and Watson (2018), Nakamura and Steinsson (2018), Bauer and Swanson (2021), Bauer and Swanson (2022), Bu et al. (2021), Swanson and Jayawickrema (2023). For example, Känzig (2021) shows that OPEC meetings have high-frequency impacts on oil prices that can be used to identify oil supply shocks; Janzen et al. (2018) show that cotton storage levels change the heteroscedasticity of precautionary demand shocks. Therefore, we believe that our approach can be used more broadly.
@@
Our inference procedure is based on the insight that both the reduced-form impulse response and covariance estimators involved in the structural identification can be represented as a mean of vector-valued stationary time series of cross-products of the $1$-step-ahead and $h$-step-ahead forecast errors. This representation allows us to employ the dependent wild bootstrap of Shao (2010) to make draws from the joint asymptotic distribution of the estimators without explicitly computing the long-run variance-covariance matrix estimator. This novel inference procedure can be used for many other identification approaches that include Cholesky (Christiano et al., 1999), sign restrictions (Gafarov et al., 2016), and external IV (Stock, 2008; Mertens and Ravn, 2013).
@@
The paper is structured as follows. Section 2 discusses the setup and the novel identification idea. Section 3 provides the estimation and inference algorithms with the corresponding theorems. Section 4 discusses the DSGE experiment. Section 5 has an empirical application to US monetary policy. Section 6 concludes. All the proofs and additional figures are provided in the appendix.
2 Identification in SVAR models
@@ 2.1 Motivating example
where observations $\eta_{t}\in\mathbb{R}^{n}$ and $\varepsilon_{t}$ is an i.i.d. zero-mean process with $\mathbb{E}\varepsilon_{t}\varepsilon_{t}^{\prime}=I_{n}$ that is referred to as structural shocks. [^5] The simultaneous equations model with lags was first proposed and studied in Haavelmo (1943); Mann and Wald (1943b). Suppose that we are interested in measuring the impact of the structural shock $\varepsilon_{1t}$ on the vector of observables $\eta_{t}$, which is given by the first column of $B=(b_{\cdot 1},\dots,b_{\cdot n})$. We can estimate $\Sigma_{\eta}\triangleq\mathbb{E}\eta_{t}\eta_{t}^{\prime}$ using the sample analog and then solve equation $\Sigma_{\eta}=BB^{\prime}$ for the first column of $B$. However, there is a continuum of solutions to this equation. Therefore, additional assumptions or data are required to uniquely select $B$. The paper aims to propose a novel way of identifying $b_{\cdot 1}$ using external instruments for the conditional heteroscedasticity process of $\varepsilon_{t}$ and to develop the inference procedure. This identification problem occurs in the more general SVAR models, which we will study later in the paper.
Before we generalize the novel identification proposal, let us consider a simple motivating example.
@@
such that $\xi_{it}$ is an i.i.d. sequence with $\mathbb{E}\xi_{t}=0$ and $\mathbb{E}\xi_{t}\xi^{\prime}_{t}=I_{2}$; the process $\sigma_{it}$ is a stationary ergodic process that is independent of $\xi_{it}$. The volatility process $\sigma^{2}_{1t}$ for the shock of interest may be correlated with some observable time series of $Z_{t}$. Suppose that this time series $Z_{t}$ only increases volatility of the first structural shock but does not change volatility for other shocks. We can then use it as an instrument for the identification of $b_{\cdot 1}$. In contrast to the conventional proxy-VAR identification, the instrument $Z_{t}$ may be independent of $\xi_{1t}$. Indeed, suppose that
$\mathbb{E}(Z_{t}-\mathbb{E}Z_{t})(\sigma^{2}_{it}-\mathbb{E}\sigma^{2}_{it})=\rho_{i},$ (3)
@@
$\eta^{2}_{1t}$ $=(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t})(b_{11}\varepsilon_{1t}+b_{12}\varepsilon_{2t}),$ (4)
@@
$\mathbb{E}\eta^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})$ $=b_{11}^{2}\mathbb{E}\varepsilon^{2}_{1t}(Z_{t}-\mathbb{E}Z_{t})=b_{11}^{2}\rho_{1},$ (6)
@@
The same logic applies if $Z_{t}$ is correlated with $\xi_{t}$ as long as $\rho_{1}\neq 0$, $\rho_{2}=0$ and $\mathbb{E}\xi_{1t}\xi_{2t}(Z_{t}-\mathbb{E}Z_{t})=0$.
@@
This assumption implies that the heteroscedasticity of the innovations for this time series must be correlated with the instrument $Z_{t}$. In the monetary shock example, one can by definition assume that the impact $b_{11}$ is nonzero for the federal funds rate ($i=1$). The instrument can only affect the conditional variance of the shock of interest and cannot impact the conditional correlation of the other structural shocks.
@@
A recent work (Lewis, 2021) showed that in the case when heteroscedasticity of a structural shock is persistent, one can use internal instruments $Z_{t}=\eta^{2}_{t-1}$. However, the approach can be applied more generally if appropriate external instruments are available, as in our empirical application.
@@
where $y_{t}$ is $n$-dimensional time series, $V\mu_{t}$ is a deterministic polynomial time trend (or time-varying drift for integrated processes) with $\mu_{t}=(1,t,t^{2},\dots,t^{k})^{\prime}$ for some $k$; $A_{1},\dots,A_{p},V$ are $(n\times n)$ parameter matrices, $\eta_{t}$ are serially uncorrelated forecast errors (innovations). We will assume that $p$ is finite and known.
@@
    $\eta_{t}$ is a martingale difference sequence (MDS), i.e. $\forall$ $t$ $\mathbb{E}[\eta_{t}|\mathcal{F}_{t-1}]=0$.
@@
    $\mathbb{E}||\eta_{t}||^{8+2\epsilon}<\infty$ and $\mathbb{E}|Z_{t}|^{4+\epsilon}<\infty$ for some $\epsilon>0$.
@@
Assumption 2.1 implies that innovations are white noise, which is required for the consistency of LS estimators in stationary VAR models. The martingale difference requirement with respect to sigma field $\mathcal{F}_{t-1}=\sigma(y_{t-1},y_{t-2},\dots)$ is only necessary for interpretation of the impulse response coefficients that will be introduced later in this section. Assumption 2.2 allows for many stationary models of conditional heteroscedasticity. [^6] See, for example, Assumption 2.1.(i-iii) in Brüggemann et al. (2016). This assumption, for example, allows for asymmetric conditional heteroscedasticity models like the ones considered in Rabemananjara and Zakoian (1993). Such models are ruled out by the popular residual wild bootstrap methods (Gonçalves and Kilian, 2004). We do not consider unconditional heteroscedasticity in innovations, since it would eliminate the need for structural identification. The 8-th moment condition imposed in our paper is sufficient for the consistency of the HAC estimators, but a weaker 4-th moment condition would be sufficient for the asymptotic normality of the LP estimators. Assumption 2.4 is standard in papers on structural identification of the shocks. Assumption 2.5 is required for asymptotic normality of $\hat{\gamma}$ and would be satisfied, for example, if $\eta_{t}$ is MDS with respect to a sigma field $\mathcal{F}_{t-1}$ that also includes $Z_{t}$ and past values of $\eta_{t}$ (for example, it is possible that the number of meetings in month $t$ is fully predictable based on the past economic conditions and the past number of meetings). This assumption is not required for the consistency of $\hat{\gamma}$ and the structural IRF.
@@
$C_{h,i,j}\triangleq\mathbb{E}(y_{t+h}|\eta_{t}=e_{j},\mathcal{F}_{t-1})-\mathbb{E}(y_{t+h}|\eta_{t}=0,\mathcal{F}_{t-1}),$ (13)
where $e_{j}\in R^{n}$ is a vector of zeros with $1$ in $j-$th position. These coefficients appear as regression coefficients in the $H$-step ahead autoregression representation of (12),
@@
Assumption 3 allows for an arbitrary number of unit roots, including complex unit roots, and cointegrating relations between $y_{t}$. It only rules out explosive roots. In practical terms, it means that one can apply our inference procedure to local projections estimated with or without a trend in levels without seasonal adjustment. It can be done without any pretests for trend stationarity, absence of cointegration, or unit roots. We allow both a polynomial trend as in Sims et al. (1990) and complex unit roots as in Tsay and Tiao (1990). Under this assumption, the least squares estimator for the coefficients in equations (12) and (14) is consistent and has well-defined distribution limits, as shown in Sims et al. (1990); Tsay and Tiao (1990) under conditionally homoscedastic MDS innovations $\eta_{t}$. In contrast to the earlier studies, we allow for more general (weak) white noise innovations with mixing by using strong laws of large numbers and functional central limit theorems for strong mixing sequences (De Jong and Davidson, 2000; Rio et al., 2017).
@@
The coefficients $C_{1},\dots,C_{H}$ can be estimated using local projection estimators of Jordà (2005). However, our empirical strategy uses conditional heteroscedasticity in the structural shocks in the dataset that has unit root time series with potential cointegration and time trends. Furthermore, structural impulse responses depend on both estimators of $(\Sigma_{\eta},C_{1},\dots,C_{H})$ and the covariances of external instruments $Z_{t}$. Existing inference methods cannot take into account all of these features together. Specifically, there are no proposals for joint inference of $\Sigma_{\eta}$ and $C_{0},\dots,C_{H}$ without pre-tests for unit roots. [^7] Brüggemann et al. (2016) provide an example where the commonly used residual wild-bootstrap is invalid for joint inference on $C_{i}$ and $\Sigma_{\eta}$ even in the stationary case. Their proposed procedure, residual block bootstrap, solves the issue for stationary VARs. Jentsch and Lunsford (2022) then extends this residual block bootstrap approach to allow external instrumental variables, also in the stationary case. Inoue and Kilian (2002); Mikusheva (2012) allow inference on individual components of $C_{h}$ in the presence of a unit root of a particular type, but impose an i.i.d. assumption on the innovations which is violated in our application.
Consider the following example that illustrates the distinction between the single coefficient and joint inference on impulse responses in the nonstationary case.